# Large Language Model LRM Large Reasoning Model

### v2

@@ D.2 LLM-Summary Configuration
In Figure˜5 we show the solve rate of different experimental configurations for LLM-Summary in addition to those of our baselines. We find that using tail turns $M>0$ improves downstream performance. Furthermore, in contrast to the 50-50 split between turns to summarize and tail turns that OpenHands Wang et al. (2025) uses, we find that summarizing more turns at once improves the solve rate. We thus proceed with $N=21,M=10$ for our main experiments.
D.3 Critic-Enhanced Summarizer
Because the LLM-Summary prompt presented in Figure˜11 is unable to consistently outperform the Observation Masking strategy on either solve rate or cost, we investigate whether we can increase the utility of the LLM call by integrating execution-free feedback Antoniades et al. (2025); Pan et al. (2024); Jain et al. (2025); Zhang et al. (2025); Shinn et al. (2023). For this, we rewrite the summary prompt from scratch. We present the resulting prompts in Figures˜12, 13 and 14 and the results of the experiment in Figure˜6.
Figure 6: Downstream task performance a randomly generated 150-sample subset of SWE-bench Verified Chowdhury et al. (2024) comparing the prompt in Appendix˜B with the joint critic-summarization prompt presented in this section. We find that simply prompting the model to include feedback in its summaries does not improve the solve rate and further increases the cost.
@@ D.3 Critic-Enhanced Summarizer
To elicit meaningful reflections, we prompt the LLM with guiding questions that it could reflect and provide insights on. These questions assess whether the agent is stuck or looping, aligned with the initial problem statement, reflect on the agent’s high-level solution approach with respect to the turns to summarize. Additionally, we provide few-shot examples to further guide the agents toward generating meaningful and actionable reflections Brown et al. (2020). Finally, as we did in the OpenHands-style prompt (Figure˜11), we provide the previous summary, or problem statement if none is available, and the turns to summarize to the model.
To our surprise, this style of prompt is not able to meaningfully outperform the prompt we used in our main experiments (Figure˜11), and also was $\approx 25\%$ more expensive per instance and resulted in $\approx 13\%$ longer trajectories. This means that a joint critic-summarization approach may actually worsen the trajectory elongation effect discussed in Section˜4.4 and further increase costs. We suggest that further work investigate when to summarize, what to summarize, and in what manner to understand the effects of various LLM-Summary approaches on agent trajectory lengths and downstream performance.
@@ D.4 Behavior of the Covered Context Management Strategies Across Turns
In Figure˜7, we show preliminary experimental results using the trajectory management strategies introduced in Section 3.1 on SWE-bench Lite-50 Badertdinov et al. (2024) with GPT-4.1-mini OpenAI et al. (2025). Here, we use a rolling window size of $M=5$, following SWE-agent Yang et al. (2024) and $N=21,M=10$ for the LLM-based approach paired with a slightly modified version of OpenHand’s prompt Wang et al. (2025) (see Appendix B). We observe that the Observation Masking strategy poses a strong baseline, consistently performing equally or better on the downstream task on SWE-bench Lite-50 Badertdinov et al. (2024) than the LLM-Summary approach despite using being much simpler.
Figure 7: Downstream task performance on SWE-bench Lite-50Badertdinov et al. (2024) across the context management strategies we cover in our study. Bars represent the mean across three experiments, the errors bars show the standard deviations. Surprisingly, $f_{RW}$ performs on par with the LLM-based strategy using more compute.
@@
To investigate why this is the case, and uncover potential scenarios in which the LLM-Summary strategy may be beneficial, we analyze the behavior of these strategies across turns in Figure˜8. The solid colored lines are the empirically observed results using SWE-agent Yang et al. (2024). For each trajectory management strategy, we run three experiments, yielding 150 trajectories total. To visualize these data, we use the micro-averaged mean for each turn. The dashed lines indicate the empirically grounded simulated behavior. To generate the data for these simulated trajectories, we compute the mean token consumption per token type across all experimental data available for the raw agent:
@@
where $T_{total}$ are the total number of turns $T$ we observed across all instances and experiments and $T_{local}$ is the number of turns of a single trajectory. We then generate a turn $T_{sim}=(\bar{r},\bar{a},\bar{o})$ using placeholder tokens. By repeatedly appending $T_{sim}$ we generate a simulated agent trajectory $\tau_{sim}=(T_{sim},\ldots,T_{sim})$ of arbitrary length. To generate the data for the simulated LLM-Summary and Observation Masking context management trajectories, we apply these strategies to $\tau_{sim}$. This allows us to study the expected behavior of these trajectory management strategies up to a large number of turns.
@@
Figure 9: Downstream task performance of a single experiment on a randomly generated 150-sample subset of SWE-bench Verified Chowdhury et al. (2024) across different context window sizes. We find that $M=10$ yields optimal performance.
@@
Appendix E Preliminary Findings using OpenHands
To probe generality across scaffolds, we ran a preliminary study with OpenHands (v0.43.0) on a 50-instance slice of SWE-bench Verified Chowdhury et al. (2024); Badertdinov et al. (2024) using Gemini 2.5 Flash Comanici et al. (2025) (no thinking), with turn limit of 250, LLM-Summary N=21, M=10, and Observation Masking M=10. We present the results in Table˜5 and Figure˜10. These results suggest that while per-instance costs are similar for masking and summarization, solve rate favored summarization on this slice. We hypothesize that scaffold-specific trajectory construction and observation preprocessing modulate the utility of summaries. For example, SWE-agent Yang et al. (2024) typically elides raw error logs and long traces from the dialog history, which may already mask failure signals; OpenHands Wang et al. (2025) retains richer logs, making summaries more informative. This motivates a controlled multi-scaffold study.
Figure 10: Preliminary results on the effectiveness versus efficiency tradeoff for context management strategies within OpenHands Wang et al. (2025) on the SWE-bench Verified-50 Badertdinov et al. (2024) subset. The plot compares solve rate (y-axis, ↑) against average cost per trajectory (x-axis, ↓) Gemini 2.5 Flash Comanici et al. (2025) without thinking. These results indicate that agent-specific design decisions may impact the efficiency and effectiveness of different context management strategies.
Table 5: Comparison of context management strategies for Gemini 2.5 Flash. We report Solve Rate (effectiveness, $\uparrow$) and Instance Cost (efficiency, $\downarrow$). We boldface the best-performing context management strategy for each metric and report change relative to the Raw Agent. All experiments use OpenHands Wang et al. (2025) on the SWE-bench Verified-50 subset Badertdinov et al. (2024).
Model Strategy Solve Rate ($\uparrow$) Instance Cost ($\downarrow$)
Gemini 2.5 Flash Raw Agent 40.0 2.39
Observation Masking 30.0(-25.0%) 1.14(-52.3%)
LLM-Summary 42.0(+5.0%) 1.22 (-49.0%)
@@ 4. Keep all sections concise and relevant
Figure 11: The LLM-Summary prompt we use in SWE-agent Yang et al. (2024) is a slightly modified version of the OpenHands LLM-Summary system prompt Wang et al. (2025). Additionally we pass the previous summary and the turns to summarize. If no previous summary is available we instead pass the task problem statement as initial context for the summary generation.

### v3

@@ D.2 LLM-Summary Configuration
In Figure˜6 we show the solve rate of different experimental configurations for LLM-Summary in addition to those of our baselines. We find that using tail turns $M>0$ improves downstream performance. Furthermore, in contrast to the 50-50 split between turns to summarize and tail turns that OpenHands [30] uses, we find that summarizing more turns at once improves the solve rate. We thus proceed with $N=21,M=10$ for our main experiments.
D.3 Critic-Enhanced LLM-Summary
A natural follow-up question is whether the LLM-Summary strategy could be improved by making the summarization process more intelligent. We explore this by enhancing our LLM-Summary strategy with execution-free feedback, a technique that has shown promise in scaling test-time compute for SE agents [22, 10, 40, 2]. In this approach, the LLM simultaneously generates a summary and critical analysis of the trajectory, incorporating both into the compressed context. This is akin to providing reflections within a single rollout, instead of across multiple rollouts [24].
Figure 7: Downstream task performance a randomly generated 150-sample subset of SWE-bench Verified [5] comparing the prompt in Appendix˜E with the joint critic-summarization prompt presented in this section. We find that simply prompting the model to include feedback in its summaries does not improve the solve rate and further increases the cost.
@@ D.3 Critic-Enhanced LLM-Summary
To elicit meaningful reflections, we prompt the LLM with guiding questions that it could reflect and provide insights on. These questions assess whether the agent is stuck or looping, aligned with the initial problem statement, reflect on the agent’s high-level solution approach with respect to the turns to summarize. Additionally, we provide few-shot examples to further guide the agents toward generating meaningful and actionable reflections [4]. Finally, as we did in the OpenHands-style prompt (Figure˜11), we provide the previous summary, or problem statement if none is available, and the turns to summarize to the model.
Testing on 150 samples from SWE-bench Verified [5] using SWE-agent [35], this critic-enhanced approach using the prompt presented in Figures˜12, 13 and 14 showed no improvement in solve rate over standard LLM-Summary. More concerning, we observed exacerbated trajectory elongation patterns, with critic-enhanced runs producing even longer trajectories than standard summarization. This is perhaps unsurprising: the critic’s reflections naturally encourage the agent to explore alternative solution paths, try additional debugging strategies, or reconsider its approach, all of which translate to more turns, thus driving cost and reducing efficiency gains.
This finding reinforces our central insight about trajectory elongation. While execution-free feedback aims to improve agent decision-making, it paradoxically increases computational costs by extending exploration. The critic’s guidance, rather than helping the agent efficiently recognize dead ends, provides additional avenues to pursue, further delaying termination. Furthermore, this increased cost, does not lead to increased downstream performance. This suggests that effective memory systems for AI agents require fundamental rethinking: simply adding more sophisticated feedback to summaries may compound rather than solve the efficiency challenges we identify.
@@ D.4 Behavior of the Covered Context Management Strategies Across Turns
In Figure˜8, we show preliminary experimental results using the trajectory management strategies introduced in Section 3.1 on SWE-bench Lite-50 [3] with GPT-4.1-mini [20]. Here, we use a rolling window size of $M=5$, following SWE-agent [35] and $N=21,M=10$ for the LLM-based approach paired with a slightly modified version of OpenHand’s prompt [30] (see Appendix E). We observe that the Observation Masking strategy poses a strong baseline, consistently performing equally or better on the downstream task on SWE-bench Lite-50 [3] than the LLM-Summary approach despite using being much simpler.
Figure 8: Downstream task performance on SWE-bench Lite-50[3] across the context management strategies we cover in our study. Bars represent the mean across three experiments, the error bars show the standard deviations. Surprisingly, $f_{RW}$ performs on par with the LLM-based strategy using more compute.
@@
To investigate why this is the case, and uncover potential scenarios in which the LLM-Summary strategy may be beneficial, we analyze the behavior of these strategies across turns in Figure˜9. The solid colored lines are the empirically observed results using SWE-agent [35]. For each trajectory management strategy, we run three experiments, yielding 150 trajectories total. To visualize these data, we use the micro-averaged mean for each turn. The dashed lines indicate the empirically grounded simulated behavior. To generate the data for these simulated trajectories, we compute the mean token consumption per token type across all experimental data available for the raw agent:
@@
where $T_{total}$ is the total number of turns $T$ we observed across all instances and experiments and $T_{local}$ is the number of turns of a single trajectory. We then generate a turn $T_{sim}=(\bar{r},\bar{a},\bar{o})$ using placeholder tokens. By repeatedly appending $T_{sim}$ we generate a simulated agent trajectory $\tau_{sim}=(T_{sim},\ldots,T_{sim})$ of arbitrary length. To generate the data for the simulated LLM-Summary and Observation Masking context management trajectories, we apply these strategies to $\tau_{sim}$. This allows us to study the expected behavior of these trajectory management strategies up to a large number of turns.
@@
Figure 10: Downstream task performance of a single experiment on a randomly generated 150-sample subset of SWE-bench Verified [5] across different context window sizes. We find that $M=10$ yields optimal performance.
@@
Appendix E LLM Summary Prompts
We share our prompt template for summary generation in Figure˜11. Compared to OpenHands [30], we remove the part of the prompt that aims to handle summary generation for tasks outside the SE domain, since our work is purely focused on the SE domain. In addition to the system prompt shown in Figure˜11, we provide a joint critic-summarization prompt in Figures˜12, 13 and 14. We discuss the effects of generating execution-free in Section˜D.3 and Section˜5.
@@ 4. Keep all sections concise and relevant
Figure 11: The LLM-Summary prompt we use in SWE-agent [35] is a slightly modified version of the OpenHands LLM-Summary system prompt [30]. Additionally we pass the previous summary and the turns to summarize. If no previous summary is available we instead pass the task problem statement as initial context for the summary generation.
@@ 4. Each reflection should identify one distinct problem and may include one
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims made in the abstract and introduction are discussed in detail in Section˜4 and Section˜5. Additionally, we extensively document our experimental setup in Section˜3 and take steps to ensure our experiments cover diverse configurations and find that our experimental results support our claims across all configurations.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We explicitly document the limitations of our work and the threats to the validity of our study in Section˜6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our work presents an extensive, empiric study of the cost-performance tradeoff of SE agent context management and thus does not contain theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In Section˜3 we comprehensively document our experimental setup, including the models used, their hyperparameters, and the benchmark we evaluate on. Furthermore, we provide details on our chosen configurations in Section˜D.4 and Appendix˜D.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release our LLM-summary implementation [^9] https://github.com/JetBrains-Research/the-complexity-trap for SWE-agent [35]. The experimental results of our main experiments can be openly accessed via HuggingFace [^10] https://huggingface.co/datasets/JetBrains-Research/the-complexity-trap.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section˜3 we comprehensively document our experimental setup, including the models used, their hyperparameters, and the benchmark we evaluate on. Furthermore, we provide details on our chosen configurations in Section˜D.4 and Appendix˜D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In Section˜3 we show the standard deviation of the solve rate, cost and context window size in our preliminary experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We present details on our infrastructure in Section˜3.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our findings pave the way toward effective and efficient LLM agents, offering an immediate way of reducing the impact of AI on our climate and environment.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We highlight the positive impact of reduced computational costs on our environment in Section˜7.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work presents an extensive evaluation regarding the efficiency of existing SE agent systems, using existing models and an existing benchmark and thus does not pose any such risk.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We credit the authors of SWE-bench [11], SWE-bench Lite-50 [3], SWE-bench Verified [5], SWE-agent [35] and OpenHands [30] at the appropriate locations in our work in Section˜2, Section˜3, and Section˜4.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release our implementation for generating LLM-summaries in SWE-agent [35] with this work in a well-documented manner.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: All of our experiments are conducted on existing benchmarks. Our work does not involve any crowdsourced labor.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: All of our experiments are conducted on existing benchmarks. Our work does not involve any crowdsourced labor.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Our work focuses on the efficiency on AI agents in SE, thus we explicitly discuss the use of specific LLMs throughout our work, for example in Section˜3 and Section˜4.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.